Which of the following four examples of terms are perfect cubes? Explain your answer to each ( why you put yes or why you put no)
125x^8
1y^12
8c^9
9b^27
1
Expert's answer
2019-01-29T10:45:07-0500

Algebraic expression is a perfect cube if it is representable in the form of a cube.

Hence 125x^8=(5x^2)^3; 1y^12=(1y^4)^3; 8c^9=(2c^3)^3 are perfect cubes, but 9b^27 is not a perfect cube because 9 is not a cube.